# Preserving feasible space variational quantum algorithm for solving the uncapacitated facility location problem

### 20 Dec 2023 v3

@@ V CONCLUSION
In conclusion, we proposed PFS-VQA for solving such UVPs, a combined ansatz that incorporates the advantages of HEA and QAOA+. To verify the high efficiency of PFS-VQA, we also applied QAOA, QAOA+, and HEA to UFLP, as well as compare the performance of these four algorithms. The Hamiltonians and the corresponding circuits are given as specified in Appendix B, C, and D respectively. We tested the performance of several instances of UFLP with 10, 14, and 22 qubits, details of the instances are given in Appendix A, finding that PFS-VQA always outperforms the other three algorithms (see Fig. 2 for details).
Specifically, for $m=2$, $n=2$, in Fig. 2 (a) and (b), PFS-VQA has a higher success probability than the other three algorithms, at least 54% higher. Meanwhile, PFS-VQA has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+. For $m=3$, $n=2$, in Fig. 2 (c) and (d), PFS-VQA has a success probability of at least 58% higher compared to the other three algorithms. Moreover, the circuit depth of PFS-VQA is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+. For $m=5$, $n=2$, in Fig. 2 (e) and (f), compared to the other three algorithms, PFS-VQA has at least five times higher success probability. And, the circuit depth of PFS-VQA is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+. For the largest scale instance, when $p=3$, our algorithm requires only about 125 iterations to converge to the optimal value, which is faster and better convergence compared to the other three algorithms (see Fig. 3 for details). In addition, we also compared the number of CNOT gates, parameter gates, and parameters of these four algorithms, and the results showed that our algorithm reduced the number of CNOT gates and parameter gates compared to QAOA and QAOA+, and both parameters and parameter gates are reduced compared to HEA (see Fig. 4 for details). The proposed algorithm is flexible as HEA can be replaced if a more efficient ansatz is available. Moreover, our algorithm is general and instructive for solving such UVPs.
@@ Acknowledgements.
This work is supported by the Beijing Natural Science Foundation (Grant No. 4222031), the National Natural Science Foundation of China (Grant Nos. 61976024, 61972048, 62272056), and the 111 Project B21049.
@@ Appendix B QAOA for UFLP
where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action. After the three key points are structured, we apply QAOA to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  5.
@@ Appendix C QAOA+ for UFLP
where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action. After the three key points are structured, we apply QAOA+ to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  6.
@@ Appendix D Hardware Efficient Ansatz for UFLP
The Hardware Efficient Ansatz (HEA) Hardware-efficient is a generic name used for ansatzes that are aimed at reducing the circuit depth needed to implement the parametric circuit when using a given quantum hardware. A major advantage of the HEA is its multifunctionality, as it can adapt to encoding symmetries HEA Bryan ; HEA Matthew and bring the relevant qubits closer to depth reduction HEA Nikolay , as well as being particularly useful for investigating Hamiltonians similar to the device’s interactions HEA Christian . Optimizing the variational parameters of the circuit is performed on classical computers, aiming to find the optimal parameters by minimizing the expected value of the phase separator Hamiltonian. For UFLP, similar to QAOA, the phase separator Hamiltonian is given in Eq. (23). For Instance 1-Instance 5, the corresponding 2-layer quantum circuit is shown in Figure  7.

### 11 Jan 2024 v4

@@ V CONCLUSION
In conclusion, we proposed VQA-PFS for solving such UVPs, a combined ansatz that incorporates the advantages of HEA and QAOA+. To verify the high efficiency of VQA-PFS, we also applied QAOA, QAOA+, and HEA to UFLP, as well as compare the performance of these four algorithms. The Hamiltonians and the corresponding circuits are given as specified in Appendix B, C, and D respectively. We tested the performance of several instances of UFLP with 10, 14, and 22 qubits, details of the instances are given in Appendix A, finding that VQA-PFS always outperforms the other three algorithms (see Fig. 3 for details).
Specifically, for $m=2$, $n=2$, in Fig. 3 (a) and (b), VQA-PFS has a higher success probability than the other three algorithms, at least 54% higher. Meanwhile, VQA-PFS has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+. For $m=3$, $n=2$, in Fig. 3 (c) and (d), VQA-PFS has a success probability of at least 58% higher compared to the other three algorithms. Moreover, the circuit depth of VQA-PFS is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+. For $m=5$, $n=2$, in Fig. 3 (e) and (f), compared to the other three algorithms, VQA-PFS has at least five times higher success probability. And, the circuit depth of VQA-PFS is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+. For the largest scale instance, when $p=3$, our algorithm requires only about 125 iterations to converge to the optimal value, which is faster and better convergence compared to the other three algorithms (see Fig. 4 for details). In addition, we also compared the number of CNOT gates, parameter gates, and parameters of these four algorithms, and the results showed that our algorithm reduced the number of CNOT gates and parameter gates compared to QAOA and QAOA+, and both parameters and parameter gates are reduced compared to HEA (see Fig. 5 for details). Our algorithm is general and instructive for solving such UVPs.
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (Grant Nos. 62371069, 62372048, 62272056), Beijing Natural Science Foundation (Grant No. 4222031) and the 111 Project B21049.
@@ Appendix B QAOA for UFLP
where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action. After the three key points are structured, we apply QAOA to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  6.
@@ Appendix C QAOA+ for UFLP
where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action. After the three key points are structured, we apply QAOA+ to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  7.
@@ Appendix D Hardware Efficient Ansatz for UFLP
The Hardware Efficient Ansatz (HEA) Hardware-efficient is a generic name used for ansatzes that are aimed at reducing the circuit depth needed to implement the parametric circuit when using a given quantum hardware. A major advantage of the HEA is its multifunctionality, as it can adapt to encoding symmetries HEA Bryan ; HEA Matthew and bring the relevant qubits closer to depth reduction HEA Nikolay , as well as being particularly useful for investigating Hamiltonians similar to the device’s interactions HEA Christian . Optimizing the variational parameters of the circuit is performed on classical computers, aiming to find the optimal parameters by minimizing the expected value of the phase separator Hamiltonian. For UFLP, similar to QAOA, the phase separator Hamiltonian is given in Eq. (23). For Instance 1-Instance 5, the corresponding 2-layer quantum circuit is shown in Figure  8.